What is the density of CO2 at 25°C and 750 Torr?

[1] 0.472 g/L [2] 1.78 g/L

[3] 1.82 g/L

[4] 2.12 g/L

[5] None of the above 


1
Expert's answer
2022-06-14T13:22:04-0400

Given:

M(CO2) = 44.01 g mol−1

P = 750 torr

T = 25°C + 273.15 = 298.15 K

R = universal gas constant = 62.3636 L torr K−1 mol−1


Solution:

Ideal gas law can be used.

Ideal gas law can be expressed as: PV = nRT

Rewrite the expression taking into account the fact that n = m / M:

MPV = mRT

Rewrite the expression taking into account the fact that d = m / V:

MP = dRT

To find the density, solve the equation for d:

d = MP / RT

d = (44.01 g mol−1 × 750 torr) / (62.3636 L torr K−1 mol−1 × 298.15 K) = 1.775 g L−1 = 1.78 g L−1

d = 1.78 g L−1


Answer: [2] 1.78 g/L